# Cambridge CB3 0WA, UK

### v2

@@ 1 Introduction
[…]
@@
where ${\cal R}$ is the scalar curvature of the induced metric on the slice $h_{ab}$. When these constraints are satisfied on a family of slices foliating the space, the gravitational action
@@
Due to the above signs, the total mass, $M$ can now be arbitrarily negative for real Euclidean-signature asymptotically-flat or asymptotically-AdS spacetimes, even when (1) and (2) are satisfied.
@@
Below, we provide various examples of families of real Euclidean configurations for which the Euclidean gravitational action is unbounded below. (We restrict throughout to connected manifolds.) Section 2 provides such examples satisfying cosmological boundary conditions, by which we mean that each slice $\Sigma$ in the given foliation is taken to be a closed manifold (without boundary). We exhibit such families for negative, positive, or zero cosmological constant. Section 3 then presents examples that are asymptotically Euclidean or asymptotically flat with thermal boundary conditions. In Section 4 we give examples that satisfy asymptotically locally AdS boundary conditions and, in particular, which are asymptotic to thermal AdS. We then turn in Section 5 to fixing the Ricci scalar instead of imposing the constraints and show that this does not provide a lower bound to the action. In Section 6 we show that, if one allows the type of complex metrics that arise from analytically continuing a real Lorentzian rotating solution, one can also easily obtain configurations with constant $R<0$ and unbounded action. We conclude with some final discussion in Section 7.
@@ 2 The case of compact slices
However, the problem also manifests itself as a purely infrared issue. To see this, it is useful to begin by considering the case where the full $D$-dimensional spacetime is a $D$-sphere of radius $r_{0}$. The Einstein-Hilbert term $-\frac{1}{16\pi}\int\sqrt{g}\,R\,{\rm d}^{D}x$ in the action is then negative. A negative cosmological constant will only make the action more negative and, in the case of a positive cosmological constant, we can still make the full action negative by taking $r_{0}$ to be sufficiently small.
@@
For spacetime dimensions $D\geq 2$, any positive contribution from the seams where two spheres are joined scales to zero with a positive power of $\epsilon$ and gives only a small correction to the action of the disjoint union. Indeed, the action remains negative and linear in the number of spheres even when the seams are smoothed out and when $\epsilon/r_{0}$ is a small number of order, say, .1.1.1.1. We thus see that large negative actions can arise due to IR effects; there is no need for curvatures to become parametrically large.
@@ 2.1 Positive cosmological constant
In this section we show that one can join together a series of spheres in a way that satisfies the constraints on each slice $\Sigma$. To do so, it is convenient to again take each $\Sigma$ to be topologically a $(D-1)$-sphere, but to allow the metric on each slice to be somewhat squashed. In particular, let us consider the case $D=4$ and take the spacetime to have the form of a so-called locally rotationally symmetric Bianchi IX cosmological model (see e.g. Misner:1973prb with $\beta_{-}=0$). We write the spacetime metric in the form
@@
where $\dot{\,}$ represents a derivative with respect to $\tau$ and we have used
@@
representing the usual round Euclidean 4-sphere. However, the behavior is more interesting for $20<\varepsilon<$ 2. Since the left-hand side of (10) is positive definite, the constraint then requires the final term on the right-hand side to be larger than the others. This condition then imposes a lower bound on $A(x)$ since the 2nd term on the right-hand side would clearly dominate in the limit $A(x)\rightarrow 0$. In addition, it also imposes an upper bound since the first term on the right clearly dominates in the limit $A(x)\rightarrow+\infty$.
Indeed, it is easy to check that the right-hand side of (10) vanishes at $A(x)=1$ for any value of $\varepsilon$. It is also straightforward to check that this defines an upper turning point, which we call $A^{+}$. For any $20<\varepsilon<$ 2, there is another turning point, $A^{-}$, which for small enough $\varepsilon$ is well approximated by
@@ 3 Asymptotically Euclidean and thermal asymptotically flat examples
Now, one might ask if the above might somehow be an artifact of taking the foliation to pinch off (so that the last leaf in the foliation is the degenerate metric with $a=0$). However, if one considers spacetimes with two disjoint asymptotically Euclidean boundaries (i.e., spacetime wormholes), it is straightforward to construct examples with nondegenerate foliations by instead choosing the curve $\gamma$ after bounce $12n-$ 1 to again follow a timelike curve from $(\alpha_{-},\beta_{-})$ that coincides with $\gamma_{0}$ at sufficiently large positive values of $\alpha$.
@@ 4 Asymptotically thermal AdS examples
In five spacetime dimensions it is also straightforward to numerically find pure vacuum examples asymptotic to thermal AdS. Using the $\sigma_{i}$ defined in (2.1), let us define the following 1--forms
@@ 5 Rescaling to constant scalar curvature
where $C$ and $\tau_{0}$ are constants of integration. For $-1<C<1$, the above solution represents a periodic configuration with period $\Delta\tau=\pi L$.
@@ 5.1 Thermal AdS4 boundary conditions
Although the potential $V(\Phi)$ is negative - becoming exponentially negative at large $\Phi$ solutions to the equations of motion derived from Eq. (49a) obey a positive energy theorem Boucher:1984yx ; Townsend:1984iu ; Amsel:2007im . The origin of this theorem lies in the fact that $V$ admits two real prepotentials, $W^{\pm}$, for all values of $\Phi$ Amsel:2007im . In particular, [^8] In Amsel:2007im , the existence of $W^{-}$ was shown to guarantee the existence of $W^{+}$, but not vice versa.
@@
With this choice of scalar profile, we can compute the on-shell action analytically as a function of $\beta$, $y_{+}$, $\gamma_{0}$, and $n$. One finds
@@ Appendix A A positivity result for linearized Euclidean Einstein-Hilbert gravity on the constraint surface
where $d$ is the dimension of the slice $\Sigma$, $\tilde{\pi}^{ij}$ is the rank-2 tensor density canonically conjugate to $h_{ij}$ and $E_{ADM}$ is the appropriate ADM boundary term and $N,N^{i}$ are the Euclidean lapse and shift and ~,~$\tilde{\cal H},\tilde{\cal H}_{i}$ are the constraints. The tildes indicate that the constraints have been written as densities on each slice $\Sigma$.
@@
[…]
@@
00 0 $=$ $\delta{\cal R}={\cal R}_{ij}\delta h^{ij}+h^{ij}\delta{\cal R}_{ij}$ (71)
@@
[…]

### v3

@@ 1 Introduction
There are suggestions in the literature that negative Euclidean actions are a result of degrees of freedom that violate the gravitational constraints. Such suggestions date back at least to the 1987 work of Hartle and Schleich Hartle:2020glw ; Schleich:1987fm which showed that, given the standard foliation of Minkowski space by constant time slices defined in some inertial frame, the Euclidean action for linearized gravity in flat Euclidean space becomes positive definite when the linearized constraints hold on each slice. Here the term constraints is used to describe equations of motion that arise in the 2nd-order formalism that restrict the allowed Cauchy data; i.e., which contain at most first derivatives normal to each slice [^1] In contrast, in the canonical formalism the off-shell action is a function of independent choices of lapse, shift, momenta, and induced metrics on each slice. In particular, at this stage no relation is imposed between the momenta and extrinsic curvatures. As a result, enforcing the associated constraints and taking the Euclidean versions of the above quantities to be real is then insufficient to make the Euclidean action bounded below, even in the linearized theory. This is analogous to the fact that the off-shell Euclidean action $S_{E}=\int d\tau(p_{E}\dot{q}+H_{E})$ is unbounded below even for a simple Harmonic oscillator (for which $H_{E}=-\frac{p_{E}^{2}}{2m}+\frac{1}{2}kq^{2}$). In constrast, imposing the equation of motion $p_{E}=\dot{q}/2m$ gives the positive-definite result $S_{E}=\frac{1}{2}\int dt_{E}(m\dot{q}^{2}+kq^{2})$.. As described in appendix A, the Hartle-Schleich argument generalizes readily to linearized gravity with a negative cosmological constant using the the global or Poincaré foliation of anti de-Sitter space. We also take the opportunity in appendix A to transcribe the argument into a form that is more transparent for our current purposes. A similar result was also established at the non-perturbative level by Hajicek Hajicek:1984nb for a toy model of Euclidean gravity. Comments in Kuchar:1970mu ; Dasgupta:2001ue ; Mazur:1989by have also been interpreted as supporting similar conclusions in Einstein-Hilbert gravity. As a result, some recent papers Banihashemi:2022htw ; Banihashemi:2024weu have taken as fact the claim that the Euclidean Einstein-Hilbert action becomes bounded below when one imposes the constraints defined by a smooth foliation.
@@
[…]
@@
where $M$ denotes the total mass of the spacetime. The positive energy theorem in general relativity shows that $M\geq 0$ if the Lorentzian constraints are satisfied. But since (1) has changed the sign of the extrinsic curvature terms, $M$ can now be arbitrarily negative even when the Euclidean constraints are satisfied.
@@
Below, we provide various examples of families of real Euclidean configurations for which the Euclidean gravitational action is unbounded below. (We restrict throughout to connected manifolds.) None of our examples will satisfy the full set of Einstein’s equations. In sections 2 - 4, the examples satisfy the Euclidean constraints, and in sections 5 and 6 they have constant Ricci scalar but do not otherwise satisfy the equations of motion. Section 2 provides examples satisfying cosmological boundary conditions, by which we mean that each slice $\Sigma$ in the given foliation is taken to be a closed manifold (without boundary). We exhibit such families for negative, positive, or zero cosmological constant. Section 3 then presents examples that are asymptotically Euclidean or asymptotically flat with thermal boundary conditions. In Section 4 we give examples that satisfy asymptotically locally AdS boundary conditions and, in particular, which are asymptotic to thermal AdS. We then turn in Section 5 to fixing the Ricci scalar instead of imposing the constraints and show that this does not provide a lower bound to the action. In Section 6 we show that, if one allows the type of complex metrics that arise from analytically continuing a real Lorentzian rotating solution, one can also easily obtain configurations with constant $R<0$ and unbounded action. We conclude with some final discussion in Section 7.
@@ 2 The case of compact slices
However, the problem also manifests itself as a purely infrared issue. To see this, it is useful to begin by considering the case where the full $D$-dimensional spacetime is a $D$-sphere of radius $r_{0}$. The Einstein-Hilbert term $\frac{1}{16\pi}\int\sqrt{g}\,R\,{\rm d}^{D}x$ in the action is then negative. A negative cosmological constant will only make the action more negative and, in the case of a positive cosmological constant, we can still make the full action negative by taking $r_{0}$ to be sufficiently small.
@@
For spacetime dimensions $D\geq 2$, any positive contribution from the seams where two spheres are joined scales to zero with a positive power of $\epsilon$ and gives only a small correction to the action of the disjoint union. Indeed, the action remains negative and linear in the number of spheres even when the seams are smoothed out and when $\epsilon/r_{0}$ is a small number of order, say, .1.1. We thus see that large negative actions can arise due to IR effects; there is no need for curvatures to become parametrically large.
@@ 2.1 Positive cosmological constant
In this section we show that one can join together a series of spheres in a way that satisfies the constraints on each slice $\Sigma$. To do so, it is convenient to again take each $\Sigma$ to be topologically a (D1)(D-1)-sphere, but to allow the metric on each slice to be somewhat squashed. In particular, let us consider the case $D=4$ and take the spacetime to have the form of a so-called locally rotationally symmetric Bianchi IX cosmological model (see e.g. Misner:1973prb with $\beta_{-}=0$). We write the spacetime metric in the form
@@
[…]
@@
representing the usual round Euclidean 4-sphere. However, the behavior is more interesting for $0<\varepsilon<2$. Since the left-hand side of (10) is positive definite, the constraint then requires the final term on the right-hand side to be larger than the others. This condition then imposes a lower bound on $A(x)$ since the 2nd term on the right-hand side would clearly dominate in the limit $A(x)\rightarrow 0$. In addition, it also imposes an upper bound since the first term on the right clearly dominates in the limit $A(x)\rightarrow+\infty$.
Indeed, it is easy to check that the right-hand side of (10) vanishes at $A(x)=1$ for any value of $\varepsilon$. It is also straightforward to check that this defines an upper turning point, which we call $A^{+}$. For any $0<\varepsilon<2$, there is another turning point, $A^{-}$, which for small enough $\varepsilon$ is well approximated by
@@ 3 Asymptotically Euclidean and thermal asymptotically flat examples
Now, one might ask if the above might somehow be an artifact of taking the foliation to pinch off (so that the last leaf in the foliation is the degenerate metric with $a=0$). However, if one considers spacetimes with two disjoint asymptotically Euclidean boundaries (i.e., spacetime wormholes), it is straightforward to construct examples with nondegenerate foliations by instead choosing the curve $\gamma$ after bounce 2n12n-1 to again follow a timelike curve from $(\alpha_{-},\beta_{-})$ that coincides with $\gamma_{0}$ at sufficiently large positive values of $\alpha$.
@@ 4 Asymptotically thermal AdS examples
In five spacetime dimensions it is also straightforward to numerically find pure vacuum examples asymptotic to thermal AdS. Using the $\sigma_{i}$ defined in (2.1), let us define the following 1-forms
@@ 5 Rescaling to constant scalar curvature
where $C$ and $\tau_{0}$ are constants of integration. For 1<C<1-1<C<1, the above solution represents a periodic configuration with period $\Delta\tau=\pi L$.
@@ 5.1 Thermal AdS4 boundary conditions
Although the potential $V(\Phi)$ is negative - becoming exponentially negative at large $\Phi$ - solutions to the equations of motion derived from Eq. (49a) obey a positive energy theorem Boucher:1984yx ; Townsend:1984iu ; Amsel:2007im . The origin of this theorem lies in the fact that $V$ can be expressed in terms of another function $W$, called the pre-potential. In fact, positive energy requires that $V$ can be recovered from two different prepotentials, $W^{\pm}$, for all values of $\Phi$ Amsel:2007im . In particular, [^8] In Amsel:2007im , the existence of $W^{-}$ was shown to guarantee the existence of $W^{+}$, but not vice versa.
@@
With this choice of scalar profile, we can compute the action analytically as a function of $\beta$, $y_{+}$, $\gamma_{0}$, and $n$. One finds
@@ Appendix A A positivity result for linearized Euclidean Einstein-Hilbert gravity on the constraint surface
where $d$ is the dimension of the slice $\Sigma$, $\tilde{\pi}^{ij}$ is the rank-2 tensor density canonically conjugate to $h_{ij}$ and $E_{ADM}$ is the appropriate ADM boundary term and $N,N^{i}$ are the Euclidean lapse and shift and $\tilde{\cal H},\tilde{\cal H}_{i}$ are the constraints. The tildes indicate that the constraints have been written as densities on each slice $\Sigma$.
@@
$0=\delta{\cal H}=\delta\left(-\sqrt{h}{\cal R}+2\sqrt{h}\Lambda\right)=\left(\delta\sqrt{h}\right){\color[rgb]{1,0,0}\cancel{\left(-{\cal R}+2\Lambda\right)}}-\sqrt{h}\delta{\cal R},$ (70)
@@
$0$ $=$ $\delta{\cal R}={\cal R}_{ij}\delta h^{ij}+h^{ij}\delta{\cal R}_{ij}$ (71)
@@
$S_{E}^{(2)}=\frac{1}{8\pi}\int_{\Sigma\times{\mathbb{R}}}\left(K_{ij}K^{ij}-{\color[rgb]{1,0,0}\cancel{K^{2}}}\right)\sqrt{h}\,{\rm d}\tau\,{\rm d}^{d}x.$ (75)